# Cosmic twins are likely rare: exploring pathways to hadron-quark stars with the same mass as neutron stars

### v1

@@ Abstract
A viable model for the dense matter equation of state above the nuclear saturation density is a hadron-to-quark phase transition at densities relevant to compact objects. In this case, stable hybrid hadron-quark stars can arise. An even more interesting scenario is one where the hadron-to-quark phase transition results in the emergence of a third branch of stable compact objects (in addition to white dwarfs and neutron stars). Inherent to the presence of a third family of compact stars is the existence of twin stars – hybrid stars with the same mass as the corresponding neutron stars, but with smaller radii. Interestingly, the neutron star-twin star scenario is consistent with GW170817. If twin stars exist in nature, it raises a question about the mechanism that leads to their formation. Here, we explore gravitational collapse as a pathway to the formation of low-mass twin stars. We perform fully general relativistic simulations of the collapse of a stellar iron core, modeled as a cold degenerate gas, to investigate whether the end product is a neutron star or a twin star. Our simulations show that even with unrealistically large perturbations in the initial conditions, the core bounces well below the hadron-to-quark phase transition density, if the initial total rest mass is in the twin star range. Following cooling, these configurations produce neutron stars. We find that twin stars can potentially form due to mass loss, e.g., through winds, from a slightly more massive hybrid star that was initially produced in the collapse of a more massive core or if the maximum NS mass is below the Chandrasekhar mass limit. The challenge in producing twin stars in gravitational collapse, in conjunction with the fine tuning required because of their narrow mass range, suggest the rarity of twin stars in nature.
@@ I Introduction
One of the most important open questions in nuclear physics research is the state of matter above the nuclear saturation density, $\rho_{s}\approx 2.7\times 10^{14}$ g cm-3. The equation of state (EOS) of cold neutron-rich matter is reasonably well understood at densities below $\sim 2\,\rho_{s}$ [1, 2, 3] (see also [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]). Uncertainties exist even for finite temperature dense nuclear matter, see e.g., [14, 15, 16]. Theoretical understanding at these densities is aided by laboratory experiments involving heavy ion collisions, see e.g., [17, 18, 19, 20, 21, 22, 23]. However, the dense-matter cold EOS remains highly uncertain at densities of $\sim 2-40\,\rho_{s}$ because of limitations in theoretical or lattice quantum chromodynamic (QCD) approaches, see, e.g., [24, 25, 26, 27, 3]. As a result, a plethora of theoretical possibilities for how matter behaves at densities higher than $2\,\rho_{s}$ is currently allowed. In this work we focus on the possibility that a quark deconfinement phase transition can take place at densities encountered inside compact objects.
@@
Figure 1: Mass-central density plot for a sequence of TOV stars based on the EOS detailed in Table 1. Stars with a fixed rest mass $M_{0}=M_{0,s}\sim M_{ch}$ are designated by labels A, B, C, D, and E. The inset zooms in on the twin star part of the curve.
If quark deconfinement takes place at astrophysically relevant densities, an interesting scenario arises when the phase transition occurs such that a “third family” of compact objects (in addition to white dwarfs and neutron stars) emerges [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Compact objects in this third family are hybrid hadron-quark stars or hybrid stars – configurations with a quark core enveloped by a hadronic shell. The stable third family of compact stars is separated from the branch of stable neutron stars (NSs) by an unstable branch of compact objects, just like stable NSs are separated from stable white dwarfs (WDs) by an unstable branch. Fig. 1 depicts these branches on a mass versus central density plot (further details on this plot are discussed in Sec. II). As a result, there exists a set of stable hybrid stars (HSs) that have the same mass as NSs but smaller radii. These HSs are typically referred to as twin stars (TSs) [54, 64, 65, 66, 67, 68, 69, 70, 71]. Recent works suggest that current experimental and multimessenger observations of NSs are either consistent with or provide evidence in favor of the existence of hadron-to-quark phase transitions in their cores [72, 73, 74, 75, 76, 77], but see [78, 79]. The third family of compact objects with potentially low-mass TSs has been shown to be consistent with a number of experiments and observations [34, 80, 81, 67, 82, 83, 84, 85]. Twin stars are particularly interesting because proof of their existence would provide valuable insights into the quark-gluon phase diagram.
The third family can arise when the EOS features a strong hadron-quark phase transition (i.e. characterized by a sufficiently large jump in energy density at a roughly fixed pressure) which is first-order (meaning that the first derivative of free energy with respect to a thermodynamic variable is discontinuous) [53]. Alternatively, hybrid EOSs with mixed pasta phases can also lead to the formation of the third family, e.g., [86]. What determines the nature of the phase transition is the surface tension between the quark and hadron phases [87, 88]. If the tension between these phases is low, a mixed phase of quark and nucleonic matter forms in-between purely nuclear and quark matter phases. By contrast, if the tension is high, a sharp transition boundary is favorable. Both possibilities are theoretically allowed, because the surface tension is not known accurately.
@@
Stellar gravitational collapse appears to be a natural pathway for the formation of HSs. One might expect this, given that WDs and NSs are endpoints of stellar evolution, see e.g., [89, 90]. However, forming TSs through gravitational collapse faces an important challenge; if one imagines the collapse of stellar core at approximately constant rest mass, following cooling, the configuration would contract and encounter first the branch of stable NSs, and then the branch of TSs, assuming further contraction. Thus, it is unclear whether there exists a natural pathway during stellar collapse that would prefer to form TSs over NSs. In fact, the work of [91] treated the evolution of configurations in the TS mass range that are initially on the unstable branch, and found that these stars naturally transition to the stable NS branch. Thus, Ref. [91] concluded that it would be challenging to form TSs through gravitational collapse because in a quasi-adiabatic, constant rest-mass contraction of a proto-NS that is on its way to forming a TS, it would have to go through the unstable branch first, which appears to favor NSs.
The primary objective of this paper is to explore whether HSs, and in particular, TSs, can form as a result of the gravitational collapse of a stellar core. To investigate this pathway toward HSs formation, we perform hydrodynamic simulations of unstable WDs in full general relativity by considering several types of initial perturbations differing in the degree of violence of the collapse. The advantage of considering a WD for our initial conditions is that it models two distinct channels to forming neutron/hybrid stars: i) the collapse of a degenerate iron core of a massive star, and ii) the accretion-induced collapse of a WD [92]. Here, we do not treat detailed microphysics or neutrino effects. Instead, we focus on treating properly relativistic gravitation, which is critical when compact objects arise.
@@ II Equations of State
There exists a broad range of possibilities for the EOS in the deep interior of an NS. Numerous studies have explored different hadronic models, quark models and hybrid hadron-quark models. For comprehensive reviews readers are referred to [27, 93, 94, 95]. Apart from microscopic, so-called realistic EOSs, the EOS can also be treated phenomenologically, e.g., by giving the pressure or sound speed as a function of rest-mass or energy density, see, e.g., [96, 97, 72, 98]. Here, we restrict our discussion to the astrophysical implications of EOSs that are based on realistic models but treated phenomenologically. The EOSs we consider give rise to a third family of stable compact stars. However, to simulate the collapse of a stellar iron core or a WD, the EOSs we construct must encompass a wide range of densities, spanning the entire range from below neutron drip to the supranuclear regime. Here we describe how we construct the phenomenological EOSs we adopt.
The EOS below the nuclear saturation density down to WD densities is reasonably well understood. Our base phenomenological EOS selected for this density regime is derived from the 6-parameter piecewise polytropic EOS introduced in [99]. For the high-density regime, the EOS needs to incorporate a phase transition, so that a stable HS branch exists. The high-density EOS that we adopt here is based on the T9 EOS of [91], which is a piecewise polytropic representation of the T9 EOS used in [36], which was in turn based on the ACS-II with $\xi=0.90$ EOS of [34]. This EOS describes zero-temperature matter in beta-equilibrium. Its hadronic part is derived from [100], the quark phase builts upon the MIT bag model [101, 102, 103, 104], and the low-density crust component is added from the models in [105, 106]. The quark phase of the original EOS adopts the constant sound speed parametrization
@@
where $\rho_{0}$ is rest-mass density, and $\rho_{0,i}\leq\rho_{0}\leq\rho_{0,{i+1}}$ is the density range of each polytropic segment. Piecewise polytropes are frequently employed in the context of the NS EOS, as many proposed tabulated, realistic nuclear EOSs can be well-fitted by them, see, e.g., [96, 7, 107, 98].
@@
After we join the low- and high-density EOSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [89] for determining general relativistic hydrostatic equilibrium stellar configurations. We then make small modifications to the EOSs to satisfy the following set of conditions:
@@
Following small modifications to the piecewise polytropic parameters of our EOS such that all aforementioned conditions are met we constructed the EOSs presented in Table 1. The EOS designated as “EOS I” in the table is used throughout most of this study. “EOS II” on the other hand is used only in one our simulations of gravitational collapse to a hybrid star above the twin star mass range. In EOS II, the maximum mass of the NS branch is slightly lower than $M_{ch}$, while the other characteristics are almost identical to those of EOS I. For the remainder of this paper, we always refer to EOS I of Table 1 unless EOS II is explicitly specified. A plot of these EOSs is provided in App. A.
@@
Table 1: The two polytropic EOSs that are adopted in this work, presented as 8-branch piecewise polytropes with 16 free parameters ($\Gamma_{i}$, $\rho_{0,i}$, $k_{1}$). Setting $M_{\odot}=1$, the first coefficient is $k_{1}=20.7$, and the values of $k_{i}$ for the other segments are determined by continuity at $\rho_{0,i}$, as given by Eq. (5).
@@
At sufficiently high densities, relevant for the most massive stellar configurations, the sound speed calculated by Eq. (3) becomes superluminal. Superluminality for ultra-dense matter with realistic EOSs is not uncommon. For example, it arises in the Akmal-Pandharipande-Ravenhall (APR) EOS [114], and efforts have been made to explain this behavior of $c_{s}$ in high densities, e.g. in [115]. However, we checked that the sound speed never becomes superluminal in our dynamical spacetime simulations.
@@
With the EOS available, we solve the TOV equations for a wide range of central rest-mass densities $\rho_{c}$ between $10^{8}-10^{16}\textrm{g cm}^{-3}$. We show the resulting mass-central density plot in Fig. 1. The plot displays the total gravitational mass $M$, also known as the Arnowitt-Deser-Misner (ADM) mass, and the total rest mass $M_{0}$ of static stars as functions of their central rest-mass density, with red and blue curves, respectively. The branches of stable ($dM/d\rho_{c}>0$) and unstable ($dM/d\rho_{c}<0$) WDs, NSs and HSs are explicitly identified based on the turning point theorem [89].
Figure 2: A 2-dimensional view (top panels) of the matter distribution inside the stable NS configuration C (left) and the HS configuration E (right). For the HS, the distinct quark core is easily distinguishable by its brighter color, indicative of the sudden increase in density. The density profile of each star is also displayed in the bottom panels as a function of the radial coordinate $r$. Red color in the background stands for the pure hadronic matter, blue represents the pure quark matter, and the narrow purple region in the plot of the HS density profile denotes the zone containing a mixture of quarks and hadrons.
@@
The largest ADM mass is attained in our EOSs is by a HS. Stars on the unstable branch that separates stable NSs and stable HSs, such as configuration D, (see inset in Fig. 1), which we will refer to as unstable TSs, are entirely in the twin star mass range. Fig. 1 can help visualize the basic question in our work. During the gravitational collapse of a stellar core/WD along a constant rest-mass path (dashed green horizontal line in Fig. 1), as the remnant cools it encounters first the stable NS branch and then the stable TS branch as it contracts quasiadiabatically. This begs the following question: if the NS branch is encountered first, then how can stable twin stars form?
We point out that during the evolution the remnant entropy changes due to heating, therefore the actual evolutionary track that we describe above is not on the plane of Fig. 1, because that corresponds to zero-entropy configurations. However, core bounce occurs above the density of the NS configuration B, and as remnant cools and contracts the evolutionary path approaches the one shown by the horizontal dashed line in Fig. 1.
Figure 3: Mass-radius relationship based on the EOS used in this work. Colors and points indicate the same quantities and configurations as in Fig. 1. Solid, dashed, and dotted curves stand for stable HSs, unstable TSs, and stable NSs branches, respectively. The green shaded area represents the mass range of TSs.
In addition to mass, the radius is another important macroscopic property of compact objects. For any given EOS, solving the TOV equations also determines the radius of each configuration. Consequently, every EOS has a unique mass-radius (M-R) relation that can be probed through multimessenger observations, and subsequently, impose constraints on the EOS of dense nuclear matter (see, e.g., [120, 121, 19, 122, 123, 27] for reviews). The M-R relation based on our EOS is displayed in Fig. 3. The three compact stars labeled in Fig. 1 are also shown in this plot with the same labels. We show these stars on the rest-mass vs radius curve, because $M_{0}$ is conserved during collapse.

### v2

# Exploring pathways to forming twin stars

@@ Abstract
A viable model for the dense matter equation of state above the nuclear saturation density includes a hadron-to-quark phase transition at densities relevant to compact objects. In this case, stable hybrid hadron-quark stars can arise. An even more interesting scenario is one where the hadron-to-quark phase transition results in the emergence of a third branch of stable compact objects (in addition to white dwarfs and neutron stars). Inherent to the presence of a third family of compact stars is the existence of twin stars – hybrid stars with the same mass as the corresponding neutron stars, but with smaller radii. Interestingly, the neutron star-twin star scenario is consistent with GW170817. If twin stars exist in nature, it raises a question about the mechanism that leads to their formation. Here, we explore gravitational collapse as a pathway to the formation of low-mass twin stars. We perform fully general relativistic simulations of the collapse of a stellar iron core, modeled as a cold degenerate gas, to investigate whether the end product is a neutron star or a twin star. Our simulations show that even with unrealistically large perturbations in the initial conditions, the core bounces well below the hadron-to-quark phase transition density, if the initial total rest mass is in the twin star range. Following cooling, these configurations produce neutron stars. We find that twin stars can potentially form due to mass loss, e.g., through winds, from a slightly more massive hybrid star that was initially produced in the collapse of a more massive core or if the maximum neutron star mass is below the Chandrasekhar mass limit. The challenge in producing twin stars in gravitational collapse, in conjunction with the fine-tuning required because of their narrow mass range, suggests the rarity of twin stars in nature.
@@ I Introduction
One of the most important open questions in nuclear physics research is the state of matter above the nuclear saturation density, $\rho_{s}\approx 2.7\times 10^{14}$ g cm-3. The equation of state (EOS) of cold neutron-rich matter is reasonably well understood at densities below $\sim 2\,\rho_{s}$ [1, 2, 3] (see also [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]). Uncertainties exist even for finite-temperature dense nuclear matter; see, e.g., [14, 15, 16]. Theoretical understanding at these densities is aided by laboratory experiments involving heavy-ion collisions; see, e.g., [17, 18, 19, 20, 21, 22, 23]. However, the dense-matter cold EOS remains highly uncertain at densities of $\sim 2-40\,\rho_{s}$ because of limitations in theoretical or lattice quantum chromodynamic (QCD) approaches; see, e.g., [24, 25, 26, 27, 3]. As a result, a plethora of theoretical possibilities for how matter behaves at densities higher than $2\,\rho_{s}$ is currently allowed. In this work we focus on the possibility that a quark deconfinement phase transition can take place at densities encountered inside compact objects.
@@
Figure 1: Mass-central density plot for a sequence of TOV stars based on the EOS detailed in Table 1. Both rest mass and gravitational (ADM) mass are illustrated. Stars with a fixed rest mass $M_{0}=M_{0,s}\sim M_{ch}$ are designated by labels A, B, C, D, and E. The inset enlarges the twin star part of the curve. More details on this plot are provided in Sec. II.
If quark deconfinement takes place at astrophysically relevant densities, an interesting scenario arises when the phase transition occurs such that a “third family” of compact objects (in addition to white dwarfs and neutron stars) emerges [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Compact objects in this third family are hybrid hadron-quark stars or hybrid stars – configurations with a quark core enveloped by a hadronic shell. The stable third family of compact stars is separated from the branch of stable neutron stars (NSs) by an unstable branch of compact objects, just like stable NSs are separated from stable white dwarfs (WDs) by an unstable branch. Fig. 1 depicts these branches on a mass versus central density plot (further details on this plot are discussed in Sec. II). As a result, there exists a set of stable hybrid stars (HSs) that have the same mass as NSs but smaller radii. These HSs are typically referred to as twin stars (TSs) [54, 64, 65, 66, 67, 68, 69, 70, 71]. A number of recent works suggest that current experimental and multimessenger observations of NSs are either consistent with or provide evidence in favor of the existence of hadron-to-quark phase transitions in their cores [72, 73, 74, 75, 76, 77], but some works do not favor this possibility [78, 79]. The third family of compact objects with potentially low-mass TSs has been shown to be consistent with a number of experiments and observations [34, 80, 81, 67, 82, 83, 84, 85]. Twin stars are particularly interesting because proof of their existence would provide valuable insights into the quark-gluon phase diagram.
The third family can arise when the EOS features a strong hadron-quark phase transition (i.e. characterized by a sufficiently large jump in energy density at a roughly fixed pressure) which is first order (meaning that the first derivative of free energy with respect to a thermodynamic variable is discontinuous) [53]. Alternatively, hybrid EOSs with mixed pasta phases can also lead to the formation of the third family; see, e.g., [86]. What determines the nature of the phase transition is the surface tension between the quark and hadron phases [87, 88]. If the tension between these phases is low, a mixed phase of quark and nucleonic matter forms in between purely nuclear and quark matter phases. By contrast, if the tension is high, a sharp transition boundary is favorable. Both possibilities are theoretically allowed, because the surface tension is not known accurately.
@@
Stellar gravitational collapse appears to be a natural pathway for the formation of HSs. One might expect this, given that WDs and NSs are end points of stellar evolution; see, e.g., [89, 90]. However, forming TSs through gravitational collapse faces an important challenge; if one imagines the collapse of the stellar core at approximately constant rest mass, following cooling, the configuration would contract and encounter first the branch of stable NSs, and then the branch of TSs, assuming further contraction. Thus, it is unclear whether there exists a natural pathway during stellar collapse that would prefer to form TSs over NSs. In fact, the work of [91] treated the evolution of configurations in the TS mass range that are initially on the unstable branch, and found that these stars naturally transition to the stable NS branch. Thus, Ref. [91] concluded that it would be challenging to form TSs through gravitational collapse because in a quasiadiabatic, constant rest-mass contraction of a proto-NS that is on its way to forming a TS, it would have to go through the unstable branch first, which appears to favor NSs.
The primary objective of this paper is to explore whether HSs and, in particular, TSs, can form as a result of the gravitational collapse of a stellar core. To investigate this pathway toward HS formation, we perform three-dimensional hydrodynamic simulations of unstable WDs in full general relativity by considering several types of initial perturbations differing in the degree of violence of the collapse. The advantage of considering a WD for our initial conditions is that it models two distinct channels to forming neutron/hybrid stars: (i) the collapse of a degenerate iron core of a massive star and (ii) the accretion-induced collapse of a WD [92]. Here, we do not treat detailed microphysics or neutrino effects. Instead, we focus on treating properly relativistic gravitation, which is critical when compact objects arise.
@@ II Equations of State
There exists a broad range of possibilities for the EOS in the deep interior of a NS. Numerous studies have explored different hadronic models, quark models and hybrid hadron-quark models. For comprehensive reviews readers are referred to [27, 93, 94, 95]. Apart from microscopic, so-called realistic EOSs, the EOS can also be treated phenomenologically, e.g., by giving the pressure or sound speed as a function of rest-mass or energy density, see, e.g., [96, 97, 72, 98]. Here, we restrict our discussion to the astrophysical implications of EOSs that are based on realistic models but treated phenomenologically. The EOSs we consider give rise to a third family of stable compact stars. However, to simulate the collapse of a stellar iron core or a WD, the EOSs we construct must encompass a wide range of densities, spanning the entire range from below neutron drip to the supranuclear regime. Here we describe how we construct the phenomenological EOSs we adopt.
The EOS below the nuclear saturation density down to WD densities is reasonably well understood. Our base phenomenological EOS selected for this density regime is derived from the six-parameter piecewise polytropic EOS introduced in [99]. For the high-density regime, the EOS needs to incorporate a phase transition, so that a stable HS branch exists. The high-density EOS that we adopt here is based on the T9 EOS of [91], which is a piecewise polytropic representation of the T9 EOS used in [36], which was, in turn, based on the ACS-II with $\xi=0.90$ EOS of [34]. This EOS describes zero-temperature matter in beta equilibrium. Its hadronic part is derived from [100], the quark phase builds upon the MIT bag model [101, 102, 103, 104], and the low-density crust component is added from the models in [105, 106]. The quark phase of the original EOS adopts the constant sound speed parametrization
@@
where $\rho_{0}$ is rest-mass density, and $\rho_{0,i}\leq\rho_{0}\leq\rho_{0,{i+1}}$ is the density range of each polytropic segment. Piecewise polytropes are frequently employed in the context of the NS EOS, as many proposed tabulated, realistic nuclear EOSs can be well fitted by them; see, e.g., [96, 7, 107, 98].
@@
After we join the low- and high-density EOSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [89] for determining general relativistic hydrostatic equilibrium stellar configurations. We then make small modifications to the EOSs by fine-tuning the free parameters of the EOS to satisfy the following set of conditions:
@@
Following small modifications to the piecewise polytropic parameters of our EOS such that all aforementioned conditions are met we constructed the EOSs presented in Table 1. The EOS designated as “EOS I” in the table is used throughout most of this study. “EOS II” on the other hand is used only in one our simulations of gravitational collapse to a hybrid star above the twin star mass range. In EOS II, the maximum mass of the NS branch is slightly lower than $M_{ch}$, while the other characteristics are almost identical to those of EOS I. For the remainder of this paper, we always refer to EOS I in Table 1 unless EOS II is explicitly specified. A plot of these EOSs is provided in App. A.
@@
Table 1: The two families of piecewise polytropic EOS that are adopted in this work, presented as eight-branch piecewise polytropes with 16 free parameters ($\Gamma_{i}$, $\rho_{0,i}$, and $k_{1}$). Setting $M_{\odot}=1$, the first coefficient is $k_{1}=20.7$, and the values of $k_{i}$ for the other segments are determined by continuity at $\rho_{0,i}$, as given by Eq. (5).
@@
At sufficiently high densities, relevant for the most massive stellar configurations, the sound speed calculated by Eq. (3) becomes superluminal. Superluminality for ultradense matter with realistic EOSs is not uncommon. For example, it arises in the Akmal-Pandharipande-Ravenhall (APR) EOS [114], and efforts have been made to explain this behavior of $c_{s}$ in high densities, e.g. in [115]. However, we checked that the sound speed never becomes superluminal in our dynamical spacetime simulations.
@@
With the EOS available, we solve the TOV equations for a wide range of central rest-mass densities $\rho_{c}$ between $10^{8}$ and $10^{16}\textrm{g cm}^{-3}$. We show the resulting mass-central density plot in Fig. 1. The plot displays the total gravitational mass $M$, also known as the Arnowitt-Deser-Misner (ADM) mass, and the total rest mass $M_{0}$ of static stars as functions of their central rest-mass density, with red and blue curves, respectively. The branches of stable ($dM/d\rho_{c}>0$) and unstable ($dM/d\rho_{c}<0$) WDs, NSs and HSs are explicitly identified based on the turning point theorem [89].
Figure 2: A two-dimensional view (top panels) of the matter distribution inside the stable NS configuration C (left) and the HS configuration E (right). For the HS, the distinct quark core is easily distinguishable by its brighter color, indicative of the sudden increase in density. The density profile of each star is also displayed in the bottom panels as a function of the radial coordinate $r$. Red color in the background stands for the pure hadronic matter, blue represents the pure quark matter, and the narrow purple region in the plot of the HS density profile denotes the zone containing a mixture of quarks and hadrons.
@@
The largest ADM mass is attained in our EOSs is by a HS. Stars on the unstable branch that separates stable NSs and stable HSs, such as configuration D (see the inset in Fig. 1), which we will refer to as unstable TSs, are entirely in the twin star mass range. Fig. 1 can help visualize the basic question in our work. During the gravitational collapse of a stellar core/WD along a constant rest-mass path (dashed green horizontal line in Fig. 1), as the remnant cools it encounters first the stable NS branch and then the stable TS branch as it contracts quasiadiabatically. This begs the following question: If the NS branch is encountered first, then how can stable twin stars form?
We point out that during the evolution the remnant entropy changes due to heating; therefore, the actual evolutionary track that we describe above is not on the plane of Fig. 1, because that corresponds to zero-entropy configurations. However, core bounce occurs above the density of the NS configuration B, and as the remnant cools and contracts the evolutionary path approaches the one shown by the horizontal dashed line in Fig. 1.
Figure 3: Mass-radius relationship based on the EOS used in this work. Colors and points indicate the same quantities and configurations as in Fig. 1. Solid, dashed, and dotted curves stand for stable HS, unstable TS, and stable NS branches, respectively. The green shaded area represents the mass range of TSs.
In addition to mass, the radius is another important macroscopic property of compact objects. For any given EOS, solving the TOV equations also determines the radius of each configuration. Consequently, every EOS has a unique mass-radius (M-R) relation that can be probed through multimessenger observations and, subsequently, impose constraints on the EOS of dense nuclear matter (see, e.g., [120, 121, 19, 122, 123, 27] for reviews). The M-R relation based on our EOS is displayed in Fig. 3. The three compact stars labeled in Fig. 1 are also shown in this plot with the same labels. We show these stars on the rest-mass versus radius curve, because $M_{0}$ is conserved during collapse.